The beam from a laser is focused with a lens, reducing the area of the beam by a factor of 2. By what factor does the amplitude of the electric field increase?
Answer #1

Amplitude of the electric field is proportional to \(\sqrt{\frac{1}{\text { Area }}}\)

If area decreases by a factor of 2 , then electric field increases by \(\sqrt{2}\)